40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 192 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  7685   Wed Nov 7 19:07:45 2012 JenneUpdateCamerasNo beam seen on external camera views

I have written some scripts which collect photos, then average them together, and subtract out an averaged background (as Rana described in elog 7678). 

I am not seeing any beam spots on any of the resulting pictures. 

 

The script to get 500 pictures is

.../scripts/general/videoscripts/videocapture50

and it's inputs are {name of camera} {folder to save in} {noBeam or withBeam}, where noBeam and withBeam indicate whether or not the PSL shutter is closed.  For the saved photos to work nicely with the Matlab script, the folder to save in should be in the format (Month_day_year/CAMERA).  So today's ITMYF pics, for example, are in Nov_7_2012/ITMYF/ .

So, you run it once with the shutter open, and once again with the shutter closed.

 

To create the new picture, open ImageBkgndSubtractor.m in the same .../scripts/general/videoscripts folder, edit the top few lines (month, day, year, camera name).  Run it, and it will read through all the pictures and supply a background-subtracted output, and save the output (as well as a version where every pixel value is multiplied by 3) in the same folder as the 500 pictures.

The pictures are all saved in

/opt/rtcds/caltech/c1/scripts/general/videoscripts/photos

so really, for my example above, it would be /opt/rtcds/caltech/c1/scripts/general/videoscripts/photos/Nov_7_2012/ITMYF/, with 2 subfolders, noBeam and withBeam, and the final pictures are saved in /opt/rtcds/caltech/c1/scripts/general/videoscripts/photos/Nov_7_2012/ITMYF/ .

 

In other, semi-unrelated news, the ITMXF camera has been not working for a while.  The bottom right quad on the test mass tv has been dark for at least a week or two.  Steve, when you have a chance (after the oplevs are all taken care of), can you see if there's something obvious that's wrong?

Here are the background subtracted photos that I've taken today:

BS_PRM_7Nov2012_SpotImage_pixelsTimes3.png

ETMXF_7Nov2012_SpotImage_pixelsTimes3.png

ETMYF_7Nov2012_SpotImage_pixelsTimes3.png

ITMYF_7Nov2012_SpotImage_pixelsTimes3.png

MC2F_7Nov2012_SpotImage_pixelsTimes3.png

MC2F is included, even though you can see the spot usually, just to prove that I'm not trying to subtract away the spot!  You just can't see it in any other picture.

  7684   Wed Nov 7 17:20:01 2012 jamieUpdateAlignmentJamie's tip tilt proposal

Quote:

Steve's elog 7682 is in response to the conversation we had at group meeting re: Jamie's proposed idea of re-purposing the active tip tilts.

What if we use the active TTs for the PR and SR folding mirrors, and use something else (like the picomotors that Steve found from the old days) for our input steering?

I think we will still need two active steering mirrors for input pointing into the OMC, after the SRC, so I think we'll still need two of the active TTs there.

My thought was about using the two active TTs that we were going to use as the input PZT replacements to instead replace the PR2/3 suspensions.  Hysteresis in PR2/3 wouldn't be an issue if we could control them.

With static input pointing, ie. leaving PZT2/3 as they are, I think we could use PRM and PR2/3 to compensate for most input pointing drift.  We might have to deal with the beam in PRC not being centered on PRM, though.

Koji's suggestion was that we could replace the PZTs with pico-motors.  This would give us all the DC input pointing control we need.

So I guess the suggestion on the table is to replace PZT1/2 with pico-motor mounts, and then replace PR2/3 with two of the active tip-tilts.  No hysteresis in the PRC, while maintaining full input pointing control.

  7683   Wed Nov 7 15:51:44 2012 JenneUpdateAlignmentJamie's tip tilt proposal

Steve's elog 7682 is in response to the conversation we had at group meeting re: Jamie's proposed idea of re-purposing the active tip tilts.

What if we use the active TTs for the PR and SR folding mirrors, and use something else (like the picomotors that Steve found from the old days) for our input steering?

  7682   Wed Nov 7 15:17:15 2012 SteveUpdateAlignmentsteering option with pico motor?
We have two ready for vacuum 1.5" mirror mounts with pico motors in our hands. 


 
Attachment 1: IMG_1795.JPG
IMG_1795.JPG
  7681   Wed Nov 7 14:59:35 2012 SteveUpdatePEMhigh particle count ALART & oplevs

Quote:

The BS camber is open only. We should close ASAP

Outside air quality is 1.7- 2.2  million particles  / cf min of 0.5 micron

 Air is still bad and the chambers are closed. Before lunch  Jamie repointed the PRM oplev. Manasa and I reset oplevs: BS and ITMX.

ETMX and ETMY are fine.

SRM and ITMY oplevs needs more work.

Attachment 1: oplevs.png
oplevs.png
  7680   Wed Nov 7 11:43:29 2012 SteveUpdatePEMhigh particle count ALART

The BS camber is open only. We should close ASAP

Outside air quality is 1.7- 2.2  million particles  / cf min of 0.5 micron

Attachment 1: highpartcount.png
highpartcount.png
  7679   Wed Nov 7 09:09:02 2012 SteveUpdateAlignmentAlignment-
PRM and SRM  OSEM LL 1.5V are they misaligned?
Attachment 1: 9amNov7w.png
9amNov7w.png
  7678   Wed Nov 7 07:11:10 2012 ranaUpdateAlignmentAlignment- POY and oplevs. photos.

The way to usually do image subtraction is to:

1) Turn off the room lights.

2) Take 500 images with no beam.

3) Use Mean averaging to get a reference image.

4) Same with the beam on.

5) Subtract the two averaged images. If that doesn't work, I guess its best to just take an image of the green beam on the mirrors using the new DSLR.

  7677   Wed Nov 7 00:10:38 2012 Jenne UpdateAlignmentAlignment- POY and oplevs. photos.
Can we have a drawing of what you did, how you confirmed your green alignment as the same as the IR (I think you had a good idea 
about the beam going to the BS...can you please write it down in detail?), and where you think the beam is clipping? Cartoon-level, 20 
to 30 minutes of work, no more. Enough to be informative, but we have other work that needs doing if we're going to put on doors 
Thursday morning (or tomorrow afternoon?).

The ETMs weren't moved today, just the beam going to the ETMs, so the oplevs there shouldn't need adjusting. Anyhow, the oplevs I'm 
more worried about are the ones which include in-vac optics at the corner, which are still on the to-do list.

So, tomorrow Steve + someone can check the vertex oplevs, while I + someone finish looking briefly at POX and POP, and at POY in 
more detail.

If at all possible, no clamping / unclamping of anything on the in-vac tables. Let's try to use things as they are if the beams are getting to 
where they need to go.  Particularly for the oplevs, I'd rather have a little bit of movement of optics on the out-of-vac tables than any 
changes happening inside.

I made a script that averages together many photos taken with the capture script that Rana found, which takes 50 pictures, one after 
another. If I average the pictures, I don't see a spot. If I add the photos together even after subtracting away a no-beam shot, the 
picture us saturated and is completely white. I'm trying to let ideas percolate in my head for how to get a useful spot. 
  7676   Tue Nov 6 18:39:16 2012 DenUpdatedigital noiseifo checking system

Matlab version of ifo digital noise estimation code is almost ready. It estimates digital noise introduced by each filter bank in each model. I'm waiting for NDS group to complete function to download online data to Matlab. Now code downloads data from the past that is not great because not all _IN1 channels are recorded and some of them are recorded at lower frequencies.

There might be some useful functions in this code for other applications as I've heard during the meetings. This is what it does

  • reads model names from the input list
  • for each model
    • finds corresponding Foton file and extracts modules with sos filters and sampling rate of the model
    • finds corresponding MDL file and makes a search for subsystems with "top_names" tag and "biquad=1" tag
    • creates _IN1 channel names using module names and subsystems with "top_names" tag
    • for each channel inside the model
      • reads filter bank parameters (which filters are ON, switches, limit, offset...)
      • downloads data
      • calculates output and estimates digital noise
      • checks that output is less them limit if it is on
      • reports if something is wrong

NDS group plans to release the function to download online data this week. Hopefully, it will be possible to download ~30 channels at a time. Code will need a few minutes of data for each channel. So it will be possible to check the whole ifo during the night.

At this point I've checked 40m using DQ channels. We have ~40 IN1_DQ channels with non-empty filter banks. These are osems channels. Digital noise is low for them.

  7675   Tue Nov 6 17:22:51 2012 Manasa, JamieUpdateAlignmentAlignment- POY and oplevs

Right now, Manasa, Jamie and Ayaka are doing some finishing touches work, checking that POY isn't clipping on OM2, the second steering mirror after the SRM, and they'll confirm that POX comes out of the chamber nicely, and that POP is also still coming out (by putting the green laser pointer back on that table, and making sure the green beam is co-aligned with the beam from PR2-PR3.  Also on the list is checking the vertex oplevs.  Steve and Manasa did some stuff with the ETM oplevs yesterday, but haven't had a chance to write about it yet.

We were trying to check POY alignment using the green laser in the reverse direction (outside vacuum to in-vac) . The green laser was installed along with a steering mirror to steer it into the ITMY chamber pointing at POY.

We found that the green laser did follow the path back into the chamber perfectly; it was clipping at the edge of POY. To align it to the center of POY (get a narrower angle of incidence at the ITMY), the green laser had to be steered in at a wider angle of incidence from the table. This is now being limited by the oplev steering optics on the table. We were not able to figure out the oplev path on the table perfectly; but we think we can find a way to move the oplev steering mirrors that are now restricting the POY alignment.

The oplev optics will be moved once we confirm with Jenne or Steve.

 

[Steve, Manasa]

We aligned the ETM oplevs yesterday. We confirmed that the oplev beam hit the ETMs. We checked for centering of the beam coming back at the oplev PDs and the QPDsums matched the values they followed before the vent.

Sadly, they have to be checked once again tomorrow because the alignment was messed up all over again yesterday.

  7674   Tue Nov 6 17:07:04 2012 jamieUpdateAlignmentAS and REFL

AS: tmp6oTENk.png

REFL: tmplamEtZ.png

  7673   Tue Nov 6 16:38:37 2012 jenne, jamie, ayaka, manasaUpdateAlignmentAlignment back under control again

We had a big alignment party early this morning, and things are back to looking good.  We have been very careful not to bump or touch tables any more than necessary.  Also, we have removed the apertures from the BS and PRM, so there are no more apertures currently left in the chambers (this is good, since we won't forget).

We started over again from the PZTs, using the PRM aperture and the freestanding aperture in front of PR2, to get the height of the beam correct.  We then moved PZTs to get the beam centered on BS, ITMY, ETMY.  We had to do a little poking of PR2 (and PR3?) to get pitch correct everywhere.

We then went to ETMX to check beam pointing, and used BS to steer the beam to the center of ETMX.  We checked that the beam was centered on ITMX.

We went through and ensured that ITMX, ITMY, PRM, SRM are all retroreflecting.  We see nice MICH fringes, and we see some fringes (although still not so nice...) when we bring PRM and SRM into alignment.

We checked the AS path (with only MICH aligned), and made sure we are centered on all of the mirrors.  This included steering a little bit on the mirrors on the OMC table, in yaw.  Initially, AS was coming out of the vacuum, but hitting the side of the black beam tube.  Now it gets nicely to the table.

For both AS and REFL, we made sure there is no clipping in the OMC chamber.

I recentered the beams for AS and REFL on their respective cameras.

IPPOS was centered on the QPD.  This involved moving the first out-of-vac steering mirror sideways a small amount, since the beam was hitting the edge of the mirror.  IPANG was aligned in-vac, and has been centered on the QPD.

Right now, Manasa, Jamie and Ayaka are doing some finishing touches work, checking that POY isn't clipping on OM2, the second steering mirror after the SRM, and they'll confirm that POX comes out of the chamber nicely, and that POP is also still coming out (by putting the green laser pointer back on that table, and making sure the green beam is co-aligned with the beam from PR2-PR3.  Also on the list is checking the vertex oplevs.  Steve and Manasa did some stuff with the ETM oplevs yesterday, but haven't had a chance to write about it yet.

  7672   Mon Nov 5 20:37:01 2012 AyakaUpdateWienerFilteringPlay with wiener filtering

 I am trying to find what limits the reduction rate with wiener filtering.
I did some calculations below:

Reduction rate estimation by microphone noise
estimate.png

When the instrumental noise (noise in microphone) and noise injected to signal after the acoustic signal is injected exist, the noise cancellation rate is limited. (I will write a short document about it later.) I assumed that there is only instrumental noise and that the other noise in PMC is below enough, and calculated the cancellation rate. The instrumental noise is modeled according to the measurement before (ELOG).
The green line is the original PMC signal, the red one is PMC residual error, and the blue one is PMC residual error estimated by the cancelling rate.
Around 30 - 80 Hz, the wiener filtering seems to be already good enough. However, I do not know what limits the cancellation rate (such as 100 - 200 Hz).

 

Filtering signals

Wfilter_notch.png

I hypothesized that the wiener filter is not good because of some peaks or other noise. So I filtered the PMC signal and mic signal to see the difference.
The red line is wiener filter with no filters, the blue one is with filters (low pass, high pass, and notch).
The wiener filter seems to get smoother but the PMC residual error did not change at all.

 

  7671   Mon Nov 5 19:38:52 2012 jamie, jenne, ayaka, denUpdateAlignmentmore alignment woes

Earlier this morning we thought things were looking pretty good.  IPPOS, IPANG, and the AS and REFL spots looked like they hadn't moved too much over the weekend.  Our plan was to do a quick check of things, check clearances, etc., tweak up the oplevs, and then close up.  This is when I made the ill-fated decisions to check the table levelling.

The BS table was slightly off so I moved one of the thick disk weights off of the other disk weight that it was sitting on, and on to the table next to it.  This seemed to improve things enough so I left it there.  ITMY didn't need any adjustment, and I move a couple smaller weights around on ITMX.  Meanwhile Jenne was adjusting the output PSL power back into it's nominal range (<100mW), and re-tweaking up the mode cleaner.

When we then looked at the vertex situation again it was far off in yaw.  This was clearly evident on PZT2, where the beam was no longer centered on the PZT2 mirror and was near the edge.  This was causing us to clip at the BS aperture.

We took some deep breaths and tried to figure out what we did that could have messed things up.

Jenne noticed that we had moved slightly on the PSL QPDs, so she adjusted the PSL output pointing to re-aquire the previous pointing, and realigned the MC.  This had a very small positive affect, but not nearly enough to compensate for whatever happened.

We spent some more time trying to track down what might have changed, but were unable to come up with anything conclusive.  We then decided to see if we could recover things by just adjusting the PZT input steering mirrors.  We couldn't; recentering at PRM, BS, ITMY, and ETMY was moving us off of PR3.

Jenne suggested we look at the spot positions on the MMT mirrors.  I had checked MMT1 and it looked ok, but we hadn't looked at MMT2.  When we checked MMT2 we noticed that we were also off in yaw.  This made us consider the possibility that the BS table had twisted, most likely when I was securing the moved mass.  Sure enough, when I manually twisted BS table, by grabbing it with my hand, very little force would cause the input beam to walk much of the way across PZT2, more than accounting for the offset.  The effect was also very clearly hysteretic as well; I could twist the table a little and it would stay in the new position.

At this point we had fucked things up enough that we realized that we're basically going to have to walk through the whole alignment procedure again, for the third time this vent.  We were able to recover the PRM retro-reflection a bit, but the tip-tilts have drifted in pitch (likely again because of the table levelling).  So we're going to have to walk through the whole procedure systematically again.

Lessons learned:  Many things are MUCH more sensitive than I had been assuming.  The tip-tilts are of course ridiculous, in that lightly touching the top or bottom of the mirror mount will move it by quite a lot in pitch.  The tables are also much more sensitive than I had realized.  In particular, tightening screws can twist the table hystereticly by milliradians, which can of course completely loose the pointing.  We need to be a lot more careful.

Assuming the table hasn't moved too much we should be able to recover the alignment by just adjusting the PZTs and tweaking the pitch of the tip-tilts.  At least that's the hope.    No more touching the table.  No more leveling.  Hopefully we can get this mostly done tomorrow morning.

  7670   Mon Nov 5 13:28:15 2012 jamieUpdateGeneral40m DCC document tree

Link to the new 40m DCC Document Tree: E1200979

  7669   Mon Nov 5 10:34:48 2012 SteveUpdateVACexisting vacuum documents

Quote:

Quote:

Quote:

Quote:

Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr.  All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open.  One of the roughing pumps was also turned on.  Other stuff we didn't notice?  Bad. 

 Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind.  All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.

CLOSE CALL on the vacuum system:

Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.

The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly  accumulated to reach ~10-50 Torr

I disconnected_ immobilized and labelled the following 6 valves:  the 4 large ion pump gate valves and VC1,  VC2  of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!

This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.

The down side of this that the computer can not identify vacuum states any longer.

This vacuum system badly needs an upgrade. I will make a list.

 While I was doing the oil change of the roughing pumps I accidentally touched the 24 V adjustment knob on the power supply.

All valve closed to default condition. I realized that the current indicator was red at 0.2A  and the voltage fluctuated from 3-13V

Increased current limiter to 0.4A and set voltage to 24V     I think this was the reason for the caos of valve switching during the VME swap.

 

 I asked Jamie to read these docs so we can do some troubleshooting while at atmosphere. We have to have the all valve closed DEFAULT condition working-tested.

I'd like to reconnect all valves after this test so the state identification software could work, including interlocks.

Posted the following documents on 40m wiki.

 

http://www.ligo.caltech.edu/~ajw/40m/40mVaccum_T000054-00.pdf
http://www.ligo.caltech.edu/~ajw/40m/40mProcedures.doc 

https://dcc.ligo.org/DocDB/0015/D000338/000/D000338-00.pdf
  7668   Mon Nov 5 09:53:35 2012 JenneUpdateAlignmentAS, REFL camera shots

Today's photos:

AS:

AS_5Nov2012.png

 REFL:

REFL_5Nov2012.png

IPANG / IPPOS trends:

IPPOS_IPANG_weekend_trend.png

 

 c.f. screen caps from Friday:

Quote:

These don't show anything too interesting, but we're including them to show where the beams are right now on the cameras, so we can compare on Monday.

 

AS:

AS_2Nov2012.png

 

REFL:

REFL_2Nov2012.png

 

 

 

  7667   Sat Nov 3 10:31:59 2012 ranaUpdateAlignmentPOP, POX, POY, IPPOS, IPANG, REFL all coming out of vac

  That's good, but I request two things:

1) Check that the REFL beam is coming from the HR surface and not the AR surface. The real REFL beam should have as much power as the Faraday output. And where does the AR surface reflection go?

2) Use frame grabber to get as many images of the spot positions on the mirrors as is reasonable. Don't endanger bumping the tables again, but take what images can be gotten by remote camera views.

  7666   Fri Nov 2 21:40:04 2012 ManasaUpdateAlignmentAS, REFL camera shots

 

 To get the camera shot of AS, Y1 mirror on the path was replaced by a 99% BS and transmitted beam was directed to the camera via a 50-50 BS (ND filters were distorting the image on the camera introducing fringes).

  7665   Fri Nov 2 20:41:53 2012 JenneUpdateAlignmentAS, REFL camera shots

These don't show anything too interesting, but we're including them to show where the beams are right now on the cameras, so we can compare on Monday.

 

AS:

AS_2Nov2012.png

 

REFL:

REFL_2Nov2012.png

 

 

  7664   Fri Nov 2 19:59:54 2012 JenneUpdateAlignmentPOP, POX, POY, IPPOS, IPANG, REFL all coming out of vac

[Evan, Jenne, Jamie]

We used the green laser pointer technique to adjust the POP steering mirrors behind PR2 to get the POP backward beam out onto the table (rather, the mirrors were adjusted so that the green laser pointer, mounted on the POX table, was co-aligned with the beam between PR2 and PR3).

We were unable (why? I feel like it wasn't so hard last time) to see the POX beam, with a camera pointed at an IR card.  We ended up just waving a lens-free CCD camera around on the POX table where we expected POX to be, found the beam, and decided that if the beam was getting to the table, that was good enough.

We then waved the camera around on the POY table, and found the POY beam on the table.  We also moved ITMY up and down in pitch, and saw that the POY beam was moving, so we were satisfied that we had the correct beam.  We should go back and do this same check with POX, although I'm pretty sure that we already have the correct beam.  But checking is good.

We confirmed that IPPOS was coming out of the chambers.  I didn't end up touching any in-vac mirrors for IPPOS, since they all looked centered, and the beam on the table was already centered on the steering mirror on the out-of-vac table.

We got IPANG out of the chamber to the ETMY table.  IPANG has, after the pickoff window, an adjustable mirror, and then a fixed mirror on the BS table.  The beam was very close to the edge, in yaw, on that fixed mirror.  Jamie unclamped it and moved it so the beam was centered, then twisted it until I got beam back down at the end, centered on the first steering mirror down there.  Then Evan and I got the beam centered on the other steering mirror on the in-vac ETMY table, and got the mirror to ~the center of the first out-of-vac steering mirror.  Then Evan adjusted the other steering optics so the beam was hitting the QPD.

We then got the real REFL beam out of the chambers.  I still don't know what that ghost/fake beam is.  Anyhow, we moved PRM around, and saw that the real REFL beam moves, while the fake one doesn't.  We adjusted the adjustable REFL steering mirror in-vac such that the real REFL beam came out to the table.  Once on the AP table, we moved the PRM around again, just to be doubly/triply sure that we had the correct beam.  We put a beam splitter (found on the SP table) after the lens in the REFL path on the AP table, and put the camera on the reflected side of that BS.  This is because, like the AS port, the beam is too dim at the normal camera spot (which for REFL is the transmission through a Y1 mirror).

Jamie has centered IPPOS and IPPANG QPDs, so we should look at the weekend trend come Monday, to see what things look like, and how they drift, if at all.

 

On Monday, we should:

* Check the alignment, and the centering of beams on all mirrors one last time

* Remove all apertures from suspended optics (I think BS and PRM may be the only two that have them at this time)

* Check oplev paths for all mirrors

* Check all pickoffs / beams that need to come out of the vacuum

* Start putting on doors

  7663   Fri Nov 2 16:09:14 2012 SteveUpdateGeneralOpen-sided mount - why

Quote:

Quote:

Quote:

Suprema- SS clear edge mirror mount 2" diameter is modified for 40m vacuum use. One left and one right handed one. It's adjustment screw housing is bronze! It is not ideal for out gassing.

It will be baked and scanned. If it passes we should use it.

We may need these to bring out some pick-off beams.

 I vote against it. We don't know about the grease inside the screw bushings - scans are not everything if adjusting the screw loosens up the grease. If we need more pick off mirrors lets just make some of the kind that we already use inside for the 2" optics.

 I think Steve had these prepared in response to my question a few days ago of how badly do we need adjustability for the POX/POY mirrors?  We already have cleaned open-sided mounts that have no adjustment screws.  So as long as the beam reflects off the ITMs horizontally (which it should), we can do yaw adjustment by twisting the whole mount. We don't need super fine yaw adjustment, we just need to get the beam out, so this is probably good enough.

We should put the POY mirror on this open-sided mount (the one without screws) some time.  Perhaps even today.

 The hex adjustment screw  is removed from the brass bushing for degreasing, cleaning and baking. Newport says the  bushing is is brass not bronze.

Brass outgassing rate is 10x higher than copper that is truth. The surface area is small so  Krytox grease is ok if it is really needed in the 40m. ( It was ok in the past and it is still there )

  7662   Fri Nov 2 14:37:36 2012 JenneUpdateGeneralOpen-sided mount - why

Quote:

Quote:

Suprema- SS clear edge mirror mount 2" diameter is modified for 40m vacuum use. One left and one right handed one. It's adjustment screw housing is bronze! It is not ideal for out gassing.

It will be baked and scanned. If it passes we should use it.

We may need these to bring out some pick-off beams.

 I vote against it. We don't know about the grease inside the screw bushings - scans are not everything if adjusting the screw loosens up the grease. If we need more pick off mirrors lets just make some of the kind that we already use inside for the 2" optics.

 I think Steve had these prepared in response to my question a few days ago of how badly do we need adjustability for the POX/POY mirrors?  We already have cleaned open-sided mounts that have no adjustment screws.  So as long as the beam reflects off the ITMs horizontally (which it should), we can do yaw adjustment by twisting the whole mount. We don't need super fine yaw adjustment, we just need to get the beam out, so this is probably good enough.

We should put the POY mirror on this open-sided mount (the one without screws) some time.  Perhaps even today.

  7661   Fri Nov 2 13:20:35 2012 MashaUpdateMachineLearningFeedback controller

Quote:

Quote:

I have uploaded to my directory a directory neural_plant. The most important file is reference_plant.c, which compiles with the command

 We would appreciate some plots. Learning curves of recurrent NN working as a plant are interesting. For harmonic oscillator your RNN should not contain any hidden layers - only 1 input and 1 output node and 2 delays at each of them. Activation function should be linear. If your code is correct, this configuration will match oscillator perfectly. The question is how much time does it take to adapt.

Does FANN support regularization? I think this will make your controller more stable. Try to use more advanced algorithms then gradient descent for adaptation. They will increase convergence speed. For example, look at fminunc function at Matlab.

Hi everyone,

I've been on break this week, so in addition to working at my lab here, I've done some NN stuff. In response to Den's response to my last post, I've included learning curve plotting capabilities, 

I've explored all of the currently documented capabilities of FANN (Fast Artificial Neural Network - it's a C library) (most likely, there are additions to the library floating around in open-source communities, but I have yet to look into those). There is extensive FANN documentation on the FANN website (http://leenissen.dk/fann/html/files/fann-h.html), but I'll cut it down to the basics here:

FANN Neural Network Architectures

standard: This creates a fully connected network, useful for small networks, as in the reference plant case 

sparse: This creates a sparsely connected network (not all of the connections between all neurons exist at all times), useful for large networks, but not useful in the reference plant case, since the number of neurons is relatively small

shortcut: This creates some connections in the network which skip over various hidden layers. Not useful in the harmonic oscillator case since there are no hidden layers. Probably won't be useful in a better-modeled referrence plant since this reduces the non-linear capabilities of the model.

FANN Training  

TRAIN_INCREMENTAL: updates the weights after every iteration, rather than after each epoch. This is faster than the other algorithms for the reference plant.

TRAIN_BATCH: updates the weights after training on the whole set. This should not be used on batches of data for the reference plant, seeing as the time history dependence of the plant is smaller than the size of the entire data set. 

TRAIN_RPROP: batch training algorithm which updates the learning parameter.

TRAIN_QUICKPROP: updates the learning parameter, and uses second derivative information, instead of just first derivative, for backpropagation. 

FANN Activation Functions

FANN offers a bunch of activation functions, including a function FANN_ELLIOT, which is essentially the "signmoid like" activation function Den and I used this summer, which runs in the order of multiplication and addition. The function parameters (steepness) can also be set.

FANN Parameters

As usual, the learning parameter can be set. While over the summer we worked with lower learning parameters, in the case of the harmonic oscillator reference plant, since the error is low after the first iteration, higher learning parameters (0.9, for example), work better. However, this is a very isolated case, and, in general, lower parameters, though convergence is slower, produce more optimal results. 

The learning momentum is another parameter that can be set - the momentum factor is a coefficient in the weight adjustment equation which allows for the difference in weights beyond the previous weight to be factored in. In the case of the reference plant, a higher learning momentum (0.9) is optimal, although in most cases, a lower learning momentum is optimal so that the learning curve doesn't oscillate terribly. 

 

FANN does not explicitly include regularization, but this can be implemented by checking the MSE  at each iteration against the MSE at the n previous iterations, where n is the regularization parameter, and stopping training if there is no significant decrease (also determined by a parameter). The error bound I specified during training was 0.0001

The best result for the reference plant was obtained using FANN_TRAIN_INCREMENTAL, a "standard" architecture, a learning rate of 0.9 (as explained above) and a learning momentum of 0.9 (these values should NOT be used for highly non-linear and more complicated systems). 

I have included plots of the learning curves - each title includes the architecture, the learning algorithm, the learning parameter, and the learning momentum if I modified it explicitly.

All of my code (and more plots!) can be found in /users/masha/neural_plant

On the whole, FANN has rather limited capabilities, especially in terms of learning algorithms, where it only has 4 (+ all of the changes one can make to parameters and rates). It is, however, much more intuitive to code with and faster han the Matlab NN library, although the later has more algorithms. I'll browse around for more open-source packages. 

Best,

Masha

standard_BATCH_0p35_ref_plant_lc.pngstandard_QUICKPROP_0p35_ref_plant_lc.pngstandard_RPROP_0p35_ref_plant_lc.pngstandard_INCREMENTAL_0p35_ref_plant_lc.png

standard_INCREMENTAL_0p9_0p9_ref_plant_lc.png

  7660   Fri Nov 2 03:28:54 2012 ranaUpdateGeneralbronze bushing for 40m vac

Quote:

Suprema- SS clear edge mirror mount 2" diameter is modified for 40m vacuum use. One left and one right handed one. It's adjustment screw housing is bronze! It is not ideal for out gassing.

It will be baked and scanned. If it passes we should use it.

We may need these to bring out some pick-off beams.

 I vote against it. We don't know about the grease inside the screw bushings - scans are not everything if adjusting the screw loosens up the grease. If we need more pick off mirrors lets just make some of the kind that we already use inside for the 2" optics.

  7659   Thu Nov 1 20:20:33 2012 Den, AyakaUpdateWienerFilteringacoustic noise in PMC

We've subtracted acoustic noise from PMC using 1 EM 172 microphone. We applied a 10 Hz high-pass filter to PMC length signal and 100,200,300:30,30 to whiten the signal.We used ~10 minutes of data at 2048 Hz as we did not see much coherence at higher frequencies.

We were able to subtract acoustic noise from PMC length in the frequency range 10-700 Hz. In the range 30-50 Hz error signal is less by a factor of 10 then target signal.

psd.pngcoh.png

  7658   Thu Nov 1 19:28:48 2012 JenneUpdateAlignmentaligned, AS beam on camera

After everyone else did the hard work, I moved the AS first-on-the-table steering mirror sideways a bit so the AS beam is on the center of the mirror, then steered the beam through the center of the lens, onto the 2" 99% BS.  I also moved the camera from it's normal place (the 1% transmitted through that BS) to the AS110 PD path, as we did last vent.  We'll need to put it back before we go back to high power.

  7657   Thu Nov 1 19:26:09 2012 jamieUpdateAlignmentaligned, this time without the crying

Jamie, Jenne, Nic, Manasa, Raji, Ayaka, Den

We basically walked through the entire alignment again, starting from the Faraday.  We weren't that far off, so we didn't have to do anything too major.  Here's basically the procedure we used:

  • Using PZT 1 and 2 we directed the beam through the PRM aperture and through an aperture in front of PR2.  We also got good retro-reflection from PRM (with PRM free-hanging).  This completely determined our input pointing, and once it was done we DID NOT TOUCH the PZT mirrors any more.
  • The beam was fortunately still centered on PR2, so we didn't touch PR2.
  • Using PR3 we direct the beam through the BS aperture, through the ITMY aperture, and to the ETMY aperture.  This was accomplished by loosening PR3 and twisting it to adjust yaw, moving it forward/backwards to adjust the beam translation, and tapping the mirror mount to affect the hysteresis to adjust pitch.  Surprisingly this worked, and we were able to get the beam cleanly through the BS and Y arm apertures.  Reclamped PR3.
  • Adjusted ITMY biases (MEDM) to get Michelson Y arm retro-reflecting to BS.
  • Adjusting BS biases (MEDM) we directed the beam through the ITMX and ETMX apertures.
  • Adjusted ITMX biases (MEDM) to get Michelson X arm retro-reflecting to BS.

At this point things were looking good and we had Michelson fringes at AS.  Time to align SRC.  This is where things went awry yesterday.  Proceeded more carefully this time:

  • Loosened SR3 to adjust yaw pointing towards SRM.  We were pretty far off at SRM, but we could get mostly there with just a little bit of adjustment of SR3.  Got beam centered in yaw on SR2.
  • Loosened and adjusted SR2 to get beam centered in yaw on SRM.
  • Once we were centered on SR3, SR2, and SRM reclamped SR2/SR3.
  • Pitch adjustment was the same stupid stupid jabbing at SR2/3 to get the hysteresis to stick at an acceptable place.**
  • Looked at retro-reflection from SRM.  We were off in yaw.  We decided to adjust SRM pointing, rather than go through some painful SR2/3 iterative adjustment.  So unclamped SRM and adjusted him slightly in yaw to get the retro-reflection at BS.

At this point we felt good that we had the full IFO aligned.  We were then able to fairly quickly get the AS beam back out on the AS table.

We took at stab at getting the REFL beam situation figured out.  We confirmed that what we thought was REFL is indeed NOT REFL, although we're still not quite sure what we're seeing.  Since it was getting late we decided to close up and take a stab at it tomorrow, possibly after removing the access connector.

The main tasks for tomorrow:

  • Find ALL pick-off beams (POX, POY, POP) and get them out of the vacuum.  We'll use Jenne's new Suresh's old green laser pointer method to deal with POP.
  • Find all OPLEV beams and make sure they're all still centered on their optics and are coming out cleanly.
  • Center IPPOS and IPANG
  • Find REFL and get it cleanly out.
  • Do a full check of everything else to make sure there is no clipping and that everything is where we expect it to be.

Then we'll be ready to close.  I don't see us putting on heavy doors tomorrow, but we should be able to get everything mostly done so that we're ready on Monday.

** Comment: I continue to have no confidence that we're going to maintain good pointing with these crappy tip-tilt folding mirrors.

 

  7656   Thu Nov 1 17:37:00 2012 SteveUpdateGeneralbronze bushing for 40m vac

Suprema- SS clear edge mirror mount 2" diameter is modified for 40m vacuum use. One left and one right handed one. It's adjustment screw housing is bronze! It is not ideal for out gassing.

It will be baked and scanned. If it passes we should use it.

We may need these to bring out some pick-off beams.

Attachment 1: IMG_1764.JPG
IMG_1764.JPG
  7655   Thu Nov 1 10:58:49 2012 jamieUpdateAlignmentprogress, then setback - trend of BS table shift

Here's a plot of the BS, PRM, and MC1 suspension shadow sensor trends over the last 24 hours.  I tried to put everything on the same Y scale:

foo.png

There definitely was some shift in the BS table that is visible in the BS and PRM that seems to be settling back now.  The MC1 is there for reference to show that it didn't really move.

  7654   Thu Nov 1 10:19:11 2012 jamieUpdateAlignmentTransmitance Measurements on LaserOptik mirror

Quote:

Quote:

...Looks like the coating is out of spec at any angle for 1064nm. E11200219-v2

The coating should have very low 1064nm p transmission at 45 degrees, which the plot seems to indicate that it does.  That's really the only part of the spec that this measurement is saying anything about.    What makes you say it's out of spec?

Ok, yes, sorry, the data itself does indicate that the transmission is way too high at 45 degrees for 1064 p.

  7653   Thu Nov 1 10:13:53 2012 jamieUpdateAlignmentTransmitance Measurements on LaserOptik mirror

Quote:

...Looks like the coating is out of spec at any angle for 1064nm. E11200219-v2

The coating should have very low 1064nm p transmission at 45 degrees, which the plot seems to indicate that it does.  That's really the only part of the spec that this measurement is saying anything about.    What makes you say it's out of spec?

  7652   Thu Nov 1 08:48:42 2012 steveUpdateAlignmentprogress, then setback - trend of BS table shift

Quote:

  But these jumps in the OSEMs are all at the level of 10-20 microns. Seems like that wouldn't be enough to account for anything; 20 microns / (pend length) ~ 50-60 microradians.

 BS table and suspensions are fine.

Attachment 1: alignmentstrugle2d.png
alignmentstrugle2d.png
Attachment 2: alignmentstrugle2da.png
alignmentstrugle2da.png
Attachment 3: befpd09252012.png
befpd09252012.png
  7651   Thu Nov 1 01:51:37 2012 ranaUpdateAlignmentprogress, then setback - trend of BS table shift

  But these jumps in the OSEMs are all at the level of 10-20 microns. Seems like that wouldn't be enough to account for anything; 20 microns / (pend length) ~ 50-60 microradians.

  7650   Wed Oct 31 22:56:41 2012 DenUpdatePEMacoustic noise

Microphone preamp box had a low-pass filter at 2kHz, Ayaka changed it to 20 kHz by replacing 100pF capacitor with a 10pF.

We've measured frequency response of the box. Signal from the microphone was split into two. One path went to the box, while another was amplified by the gain 20 (and bandpass filter 1Hz - 300kHz) and sent to spectrum analyzer. Coherence and frequency response were measured using box output and amplified input. Low-pass filter in the box does not limit our sensitivity.

Acoustic noise significantly decreases at frequencies higher then 2kHz. So we need to modify the circuit by adding whitening filter.

pre_after_mic.png  freq_resp.png

I've plugged in PMC length channel into PEM board CH15 through and amplifier (gain=200) that is AC coupled to avoid ~2.5 DC V coming from PMC servo.  I measured coherence with microphone that was located ~30 cm higher. Measurements show contribution of acoustic noise to PMC length in the frequency range 20-50 Hz. In this range PMC length / MC length coherence is ~0.5.

Acoustic noise couples to PMC length in a non-stationary way. 5 minutes after the first measurement I already see much higher contribution. This was already discussed here. I've made C1:X02-MADC3_TP_CH15 a DQ channel at 64kHz. This a fast PMC length channel.

Next step will be to use several microphones located around PMC for acoustic noise cancellation.

DSC_4792.JPG    DSC_4793.JPG

Attachment 3: pmc.pdf
pmc.pdf pmc.pdf
Attachment 4: pmc_high.pdf
pmc_high.pdf
  7649   Wed Oct 31 17:36:39 2012 jamieUpdateAlignmentprogress, then setback - trend of BS table shift

Quote:

Here is a two hour set of second trends of 2 sensors per mirror, for BS, PRM, ITMY and MC1.  You can see about an hour ago there was a big change in the BS and PRM suspensions, but not in the ITMY and MC1 suspensions.  This corresponds as best we can tell with the time that Jamie was figuring out and then fixing PZT2's mount.  You can see that the table takes some time to relax back to it's original position.  Also, interestingly, after we put the doors on ~10 or 20 minutes ago, things change a little bit on all tables. This is a little disconcerting, although it's not a huge change.

 what's going on with those jumps on MC1?  It's smaller, but noticeable, and looks like around the same time.    Did the MC table jump as well?

more looking tomorrow.

  7648   Wed Oct 31 17:33:39 2012 KojiUpdateAlignmentTransmitance Measurements on LaserOptik mirror

...Looks like the coating is out of spec at any angle for 1064nm. E11200219-v2

  7647   Wed Oct 31 17:18:34 2012 JenneUpdateAlignmentprogress, then setback - trend of BS table shift

Here is a two hour set of second trends of 2 sensors per mirror, for BS, PRM, ITMY and MC1.  You can see about an hour ago there was a big change in the BS and PRM suspensions, but not in the ITMY and MC1 suspensions.  This corresponds as best we can tell with the time that Jamie was figuring out and then fixing PZT2's mount.  You can see that the table takes some time to relax back to it's original position.  Also, interestingly, after we put the doors on ~10 or 20 minutes ago, things change a little bit on all tables. This is a little disconcerting, although it's not a huge change.

Attachment 1: PRM_BS_table_bumped_ITMY_MC1_no_big_change_2hoursBack.png
PRM_BS_table_bumped_ITMY_MC1_no_big_change_2hoursBack.png
  7646   Wed Oct 31 17:11:40 2012 jamieUpdateAlignmentprogress, then setback

jamie, nic, jenne, den, raji, manasa

We were doing pretty well with alignment, until I apparently fucked things up.

We were approaching the arm alignment on two fronts, looking for retro-reflection from both the ITMs and the ETMs.

Nic and Raji were looking for the reflected beam off of ETMY, at the ETMY chamber.  We put an AWG sine excitation into ETMY pitch and yaw.  Nic eventually found the reflected beam, and they adjusted ETMY for retro-reflection.

Meanwhile, Jenne and I adjusted ITMY to get the MICH Y arm beam retro-reflecting to BS.

Jenne and I then moved to the X arm.  We adjusted BS to center on ITMX, then we moved to ETMX to center the beam there.  We didn't both looking for the ETMX reflected beam.  We then went back to BS and adjusted ITMX to get the MICH X arm beam retro-reflected to the BS.

At this point we were fairly confident that we had the PRC, MICH, and X and Y arm alignment ok.

We then moved on the signal recycling cavity.  Having removed and reinstalled the SRC tip-tilts, and realigning everything else, they were not in the correct spot.  The beam was off-center in yaw on SR3, and the SR3 reflected beam was hitting low and to the right on SR2.  I went to loosen SR3 so that I could adjust it's position and yaw, and that when things went wrong.

Apparently I hit something BS table and completely lost the input pointing.  I was completely perplexed until I found that the PZT2 mount looked strange.  The upper adjustment screw appeared to have no range.  Looking closer I realized that we somehow lost the gimble ball between the screw and the mount.  Apparently I somehow hit PZT2 hard enough to separate from the mirror mount from the frame which caused the gimble ball to drop out.  The gimble ball probably got lost in a table hole, so we found a similar mount from which we stole a replacement ball.

However, after putting PZT2 back together things didn't come back to the right place.  We were somehow high going through PRM, so we couldn't retro-reflect from ITMY without completely clipping on the PRM/BS apertures.  wtf.

Jenne looked at some trends and we saw a big jump in the BS/PRM osems.  Clearly I must have hit the table/PZT2 pretty hard, enough to actually kick the table.  I'm completely perplexed how I could have hit it so hard and not really realized it.

Anyway, we stopped at this point, to keep me from punching a hole in the wall.  We will re-asses the situation in the morning.  Hopefully the BS table will have relaxed back to it's original position by then.

  7645   Wed Oct 31 14:31:34 2012 SteveUpdatePEMhigh particle count

High particle count confirmed with #2 counter

Attachment 1: highparticlec.png
highparticlec.png
  7644   Wed Oct 31 12:58:17 2012 RajiUpdateAlignmentTransmitance Measurements on LaserOptik mirror

Quote:

I measured the transmitted power @1064nm on one of the LaserOptik mirrors labled SN6

Here is the data

Polarization Input Angle Input Power(mW) Output Power(mW) Transmittance (%)
p 0 6.2 2.67 48
p 0 100 52 52
p 45 6.2 0.76 12
p 45 100 1,5 1
s 0 8.2 3.15 38
s 0 100 40 0.4
s 45 8.2 0.5 6
s 45 100 0.66 0.006

The mirror is not a good reflector at 0 deg.

 More data on the transmission. Measured the tranmission as a funtion of incidence angle at 1064nm

Attachment 1: Transmission-plot@1064nm.pdf
Transmission-plot@1064nm.pdf
Attachment 2: Transmission-data@1064nm.pdf
Transmission-data@1064nm.pdf
  7643   Wed Oct 31 01:06:31 2012 DenUpdateAlignmentYarm

 

 Jenne, Den

We looked at beam spots on ITMY and ETMY. We switched to smaller apertures on the other side of the rulers. For ITMY beam spot was 1mm below and 1mm south (right if you look in the direction ITMY -> ETMY) from the aperture center, for ETMY - 4 mm up and 3mm north from the aperture center. We made a correction for this using PZT 1 and 2. Now beam spots are in the middle of the apertures on ITMY and ETMY.

We tried to look at reflected beam from ETMY but it was hard to see the dependence between ETMY DC offset and reflected beam. We'll continue tomorrow.

  7642   Tue Oct 30 11:51:45 2012 JenneUpdateAlignmentPRC aligned, Yarm almost aligned

[Raji, Jenne]

We tweaked PZT2, PZT1 (yaw only), and PR3 (pitch only) to get the beam ~centered on the BS aperture, the ITMY aperture, and the ETMY aperture.

After lunch I'll tweak up the MC alignment, since, although the spots are in the right places, the transmitted beam could be higher power.  This will make it easier to check our pointing, especially since the ETMY spot is larger than our aperture, but the beam is dim.

We're getting there!

  7641   Mon Oct 29 18:50:02 2012 JenneUpdateAlignmentPRC aligned, Yarm almost aligned

[Jamie, Jenne, Raji, with consultation from Nic, Ayaka and Manasa]

We went back and re-looked at the input alignment, and now we're "satisfied for the moment" (quote from Jamie) with the PRC alignment.  Also, by adjusting the PR folding mirrors, we are almost perfectly aligned to the Yarm.

What we did:

Set PRM DC biases to 0 for both pitch and yaw.

Aperture was attached to PRM cage, double aperture was attached to BS cage, free-standing aperture was placed in front of PR2. 

Adjusted PZT1, PZT2 such that we were centered on PZT2, and through apertures at PRM and PR2.   This was mainly for setting beam height in PRC.

Checked centering on PZT1, MMT1, MMT2, PZT2.

Adjusted PRM pitch bias and PZT2 yaw such that REFL beam was retro-reflected from PRM.

Checked that REFL beam came nicely out of Faraday.

Checked that beam was still going through center of PRM aperture, and pitch height at PR2 was good.

Moved PR2 sideways until beam hit center in yaw of PR2.

Twisted PR2 such that beam was hitting center of PR3.

Moved and twisted PR3 (many times) so that beam went through BS input and output apertures, and through center of ITMY aperture.

Found that beam was just getting through black glass aperture at ETMY, top left corner, if looking at the face of ETM from ITM.

Locked down dog clamps on PR2.

This required some re-adjustment of PR3.  Re-did making sure going through BS apertures and ITMY aperture, locked down PR3 dog clamps.

Found that we are centered in yaw at ETMY, a little high in pitch on ETMY.

Replaced all of the light doors, to take a break.  4 hours in bunny suits seemed like enough that we earned a break.

This all sounds more straighforward than it was.  There was a lot of iteration, but we finally got to a state that we were relatively happy with.

 

What we will do:

Tweak PZT2 a *tiny* bit in pitch, ~0.5 mrad, so that the beam goes through the ETMY aperture.

See if we can align EMTY and ITMY to get multiple bounces through the Yarm.

Remove ETMX heavy door, steer BS such that we're getting through the center of an aperture at ETMX.

Align ETMX and ITMX such that we get multiple bounces through the Xarm.

Check SRM, AS path alignment.

Check REFL out of vac alignment.

Check other pickoffs.

Check all oplevs.

Check IPPOS/IPANG

 

We have a open-sided 2" mirror mount that we are considering using for the POY pick-off mirror.  This might help us get a little more clearance in the Y-arm of the Michelson.  Problem is the mount is not steerable, so we need to determine if that's doable or not.

 

  7640   Mon Oct 29 18:14:55 2012 DenUpdateSUS PITCH damping needed

 

 We've received all parts that we need for eddy current damping. I've made an estimate of Q with dirty tip-tilt. It looks fine (Q~1)

We need to check ring magnets for vacuum compatibility. Bob start baking on Friday.

DSC_4787.JPG     DSC_4791.JPG

  7639   Mon Oct 29 14:57:41 2012 janoschUpdateGeneraltip-tilt phase maps

Quote:

 [Jan, Manasa]

Below are phasemaps for the tip-tilts with both tilt and RoC removed. We have not used Koji's code; but tweaked the earlier code to remove curvature. 

 The posted residual phase maps show circular contours since the data came with relatively low resolution in height. This is ok for what we want to do with these phase maps (i.e. simulating higher-order mode content in the PRC using Finesse). Better resolution is only required if you want to understand in detail optical scattering out of the cavity. Anyhow, the circular artifacts can be removed by first interpolating the phase maps to a higher lateral resolution, and then performing tilt and curvature subtraction. So we will soon have better looking phase maps posted. Then we should think about what type of Finesse simulation we could run. Certainly one simulation is to look at the beam shape in the PRC, but more interesting could be how sensitive the shape is to mirror alignments. The current simulation shows a mode that resembles the TEM01, but I have not yet tried to find optimal alignment of the mirrors (in simulation) to search for the TEM00 mode.

  7638   Mon Oct 29 11:27:42 2012 ManasaUpdateGeneraltip-tilt phase maps

 [Jan, Manasa]

Below are phasemaps for the tip-tilts with both tilt and RoC removed. We have not used Koji's code; but tweaked the earlier code to remove curvature.

The RoC values matched approximately to that quoted by Gari Lynn ~700m.

RoC of tip-tilts
Mirror
RoC (m)
SN1 748.7176
SN2 692.7408
SN3 707.0336
SN4 625.5152
SN5 672.5340
SN6 663.7791

 

Phasemaps

The color scale for height are not the same for all mirrors.

 

SN1, SN2 and SN3

sn1_UC_UT.pngsn2_UC_UT.pngsn3_UC_UT.png

SN4, SN5 and SN6

sn4_UC_UT.pngsn5_UC_UT.pngsn6_UC_UT.png

  7637   Mon Oct 29 09:33:42 2012 SteveUpdateSUSPRM & ETMY sus damping restored
  7636   Mon Oct 29 08:41:22 2012 AyakaUpdateAdaptive FilteringMicrophone noise again

Quote:

Quote:

The circuit noise improves so much, but many line noises appeared.
Where do these lines (40, 80, 200 Hz...) come from?
These does not change if we changed the microphones...

Anyway, I have to change the circuit (because of the low-pass filter). I can check if the circuit I will remake will give some effects on these lines.

I do not think that 1U rack power supply influenced on the preamp noise level as there is a 12 V regulator inside. Lines that you see might be just acoustic noise produced by cpu fans. Usually, they rotate at ~2500-3000 rpm => frequency is ~40-50 Hz + harmonics. Microphones should be in an isolation box to minimize noise coming from the rack. This test was already done before and described here

I think we need to build a new box for many channels (32, for example, to match adc). The question is how many microphones do we need to locate around one stack to subtract acoustic noise. Once we know this number, we group microphones, use 1 cable with many twisted pairs for a group and suspend them in an organized way.

 I do not think they are acoustic sounds. If so, there should be coherence between three microphones because I placed three at the same place, tied together. However, there are no coherence at lines between them.

ELOG V3.1.3-